ganeshie8 (ganeshie8):

Suppose \(n|(a^s-1)\) and \(n|(a^t-1)\) Can we say anything about the relationship between \(s\) and \(t\) ?

OpenStudy (anonymous):

cant see anything :\ first i thought n =a-1 but then ok not this form... \(\Large a^s\equiv 1 \mod n \\\Large a^t\equiv 1 \mod n \\ \rightarrow\Large a^{s+t}\equiv 1 \mod n \\ \text{however this seems leads to nothing :\ } \)

OpenStudy (anonymous):

i only can say W.L.O.G s>t or so lol

ganeshie8 (ganeshie8):

that also gives \[\large a^{s-t} \equiv 1 \pmod{n}\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

for that we need ind thingy

ganeshie8 (ganeshie8):

because \(\large a^s \equiv a^t \pmod{n}\) divide \(\large a^t\) both sides assuming \(\gcd(a,n)=1\)

OpenStudy (anonymous):

oh got what your saying, but this goes for strict conditions also we need to be careful \(a^{s-t}>n\) else its false

OpenStudy (anonymous):

so i won't say that applies :|

ganeshie8 (ganeshie8):

ok

OpenStudy (anonymous):

but if n=a-1 then we can say gcd(s,t)= min(s,t)

OpenStudy (anonymous):

eh nvm seems im not thinking good of it

ganeshie8 (ganeshie8):

can we say this \[\large n|(a^s-1)~~\text{and}~~n|(a^t-1) ~~\implies n|(a^{\gcd(s,t)}-1)\] ?

ganeshie8 (ganeshie8):

i dont know if thats true, but it is working for all the examplies i tried..

OpenStudy (anonymous):

its true for n=a-1

OpenStudy (anonymous):

lets see if its not

OpenStudy (anonymous):

im not getting ideas will try ltr :\ \(\Large a^s\equiv 1 \mod n \\\Large a^t\equiv 1 \mod n \\ \) assume gcd(s,t)=d s=df t=dg \(\Large a^{df}\equiv 1 \mod n \\\Large a^{dg}\equiv 1 \mod n \\ \)